# Discrete quantum groups and their duals

### v1

@@ Acknowledgments
  • Proof: i) We first look at the setting in [VD-si1]. For both the algebras $B$ and $C$ in that paper, we simply take $A$, and for $E$ we take $\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Here we have -algebras and $\Delta(h)$ is self-adjoint. We know that $\Delta(h)(1\otimes a)$ and $(a\otimes 1)\Delta(h)$ both belong to $A\otimes A$ and similarly on the other side. This shows that Assumption 1.1 of [VD-si1] are fulfilled. In particular $\Delta(h)$ is regular in the sense of that assumption.
@@
    that $S(e_{pj})\in V_{\overline{\alpha}}$ for all $j,p$. This implies that $V_{\overline{\alpha}}=A_{\overline{\alpha}}$ for all $\alpha$ and hence $V=A$. By the way, from Proposition 1.3 in [VD-si1] we can conclude that the legs of $\Delta(h)$ are all of $A$, now in the sense of our paper here. Also Proposition 1.4 in [VD-si1] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [VD-si1]. Since we know here already that
@@
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [VD-si1] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [VD-si1], we get here
@@
Recall from Notation 2.10 the definition of the element $q_{\alpha}$ in $A_{\alpha}$. We have $q_{\alpha}=\\ (\omega_{\overline{\alpha}}\otimes\iota)\Delta(h)$ where $\omega_{\overline{\alpha}}$ is the usual trace on the component $A_{\overline{\alpha}}$. It is the unique element in $A_{\alpha}$ satisfying $aq_{\alpha}=q_{\alpha}S^{2}(a)$ for all $a\in A_{\alpha}$ and normalized such that $\omega_{\alpha}(q_{\alpha})=1$. It is a strictly positive element in $A_{\alpha}$.
We then get the following formula for the left integral on $A$.
@@
  • Theorem Let $\varphi$ be the left integral on $A$, normalized such that $\varphi(h)=1$. Then for each $\alpha$ we have $\varphi(a)=\omega_{\alpha}(aq_{\alpha}^{-1})$ when $a\in A_{\alpha}$.
  • Proof: Let $\alpha$ be an index in $I$ and consider the associated data we have used before. We found that
    $\sum_{i,j}\lambda_{i}S(e_{ij})\otimes e_{ji}=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})=\Delta(h)(1_{\overline{\alpha}}\otimes 1)$
    where $(e_{ij})$ are matrix units in $A_{\alpha}$, satisfying $q_{\alpha}=\sum_{i}\lambda_{i}e_{ii}$. Then
    $((\iota\otimes\omega_{\alpha}(\,\cdot\,q_{\alpha}^{-1}))\Delta(h))$ $=\sum_{i,j}\lambda_{i}S(e_{ij})\omega_{\alpha}(e_{ji}\lambda_{i}^{-1})$
@@
    $=S(1_{\alpha})=1_{\overline{\alpha}}.$
    This proves that $\omega_{\alpha}(aq_{\alpha}^{-1})=\varphi(a)$ for $a\in A_{\alpha}$.
We now obtain formulas for the other data, the modular automorphism $\sigma$, the modular element $\delta$ and the right integral $\psi$. We also verify some formulas relating these data.
@@
  • Proposition The $\alpha$-component of the modular element $\delta$ is $\omega_{\alpha}(q_{\alpha}^{-1})\,q_{\alpha}^{2}$.
  • Proof: Take an index $\alpha$ and consider the element $q_{\alpha}\in A_{\alpha}$ as defined in Notation 2.10. Then $aq_{\alpha}=q_{\alpha}S^{2}(a)$ and $S(q_{\alpha})S(a)=S^{3}(a)S(q_{\alpha})$ for all $a\in A_{\alpha}$. As $S(A_{\alpha})=A_{\overline{\alpha}}$ we have $aS(q_{\alpha}^{-1})=S(q_{\alpha}^{-1})S^{2}(a)$ for all $a\in A_{\overline{\alpha}}$. Therefore
    $q_{\overline{\alpha}}=\frac{1}{\omega_{\overline{\alpha}}(S(q_{\alpha}^{-1}))}S(q_{\alpha}^{-1})=\frac{1}{\omega_{\alpha}(q_{\alpha}^{-1})}S(q_{\alpha}^{-1}).$
    Then for $a\in A_{\overline{\alpha}}$,
    $\varphi(a)$ $=\omega_{\overline{\alpha}}(aq_{\overline{\alpha}}^{-1})=\omega_{\overline{\alpha}}(q_{\overline{\alpha}}^{-1}a)$
    $=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(S(q_{\alpha})a)$
    and we get
    $(\varphi\otimes\iota)\Delta(h)(1\otimes 1_{\alpha})$ $=\sum_{(h)}\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(S(q_{\alpha})h_{(1)})h_{(2)}$
    $=\sum_{(h)}\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\overline{\alpha}}(h_{(1)})q_{\alpha}h_{(2)}$
    $=\omega_{\alpha}(q_{\alpha}^{-1})\,q_{\alpha}^{2}.$
@@
It is now easy to obtain the right integral $\psi$ because we have $\psi(a)=\varphi(S(a))$ since we have normalized such that $\psi(h)=\varphi(h)=1$. With this choice we get $\psi(a)=\varphi(a\delta)$ and so, for $a\in A_{\alpha}$ we find
$\psi(a)=\varphi(a\delta)=\omega_{\alpha}(a\delta q_{\alpha}^{-1})=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(aq_{\alpha}).$
@@
$\varphi(ab)=\omega_{\alpha}(abq_{\alpha}^{-1})=\omega_{\alpha}(bq_{\alpha}^{-1}a)=\varphi(bq_{\alpha}^{-1}aq_{\alpha}).$
We see that $aq_{\alpha}=q_{\alpha}\sigma(a)$ and this is in agreement with $\sigma(a)=S^{2}(a)$.
Similarly, again for $a,b\in A_{\alpha}$,
$\psi(ab)=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(abq_{\alpha})=\omega_{\alpha}(q_{\alpha}^{-1})\omega_{\alpha}(bq_{\alpha}a)=\psi(bq_{\alpha}aq_{\alpha}^{-1}).$
We see that $\sigma^{\prime}(a)=q_{\alpha}aq_{\alpha}^{-1}$ and so $aq_{\alpha}^{-1}=\sigma^{\prime}(a)q_{\alpha}^{-1}$. This is in agreement with $\sigma^{\prime}(a)=S^{-2}(a).$
@@
In this section, we will consider the dual of a discrete quantum group. We will follow the approach to duality as developed for general algebraic quantum groups in [VD-alg]. The dual will be a compact quantum group and we know that all compact quantum groups arise in this way. By looking at a compact quantum group as the dual of a discrete quantum group, we can obtain results about compact quantum groups in a different and easier way. We refer to the appendix for more comments on this matter.
@@
    iii) In this case, we also use $\Delta$ for the coproduct, $\varepsilon$ for the counit, $S$ for the antipode, etc. on $B$.
@@
Observe also the presence of the inverse of the antipode in Equation (3.3). This is one of the consequences of this convention. For the same reason, there is the lack of symmetry between (3.4) and (3.5). We do not get the usual formulas for a pairing of Hopf -algebras or multiplier Hopf -algebras (as e.g. in [Dr-VD]), but rather a modified form.
@@
So we see that the dual of the discrete quantum group (seen as a multiplier Hopf -algebra with positive integrals) is a Hopf -algebra with positive integrals. We also see from the above equation that the integral is faithful. This is in fact automatic (see e.g. [VD-alg]). We will encounter a similar result with a similar argument in Proposition 3.13 later in this section, as well in Proposition 4.6 the next section with a dual form.
@@
  • Proof: i) Let $d=\varphi(a\,\cdot\,)$ with $a\in A$. Then for all $c\in B$,
@@
    $=\sum_{(h)}\langle h_{(1)},c\rangle\varphi(ah_{(2)})$
    $=\sum_{(h)}\langle S(a)h_{(1)},c\rangle\varphi(h_{(2)})$
@@
    $=\sum_{(h)}\varphi(ah_{(1)})\langle h_{(2)},b\rangle)$
    $=\sum_{(h)}\varphi(h_{(1)})\langle S^{-1}(a)h_{(2)},b\rangle)$
@@
Formulas of this type are found at various places in the literature. One possible reference is [De-VD1] where they are obtained in the more general case of algebraic quantum hypergroups. They are also found in [VD-part1].
If we let $B$ act on the GNS-representation space associated with this integral, we can complete it to a C-algebra. Also the coproduct can be extended. One of the key features that makes this possible is the fact that the -algebra is represented by bounded operators in the GNS-space associated with the positive integral. This is not true for general positive linear functionals on a -algebra. This result is already found in [Ku-VD], see also [DC-VD] and the more recent approach in [VD-part2].
In what follows we will sometimes denote the integral on the dual compact quantum group $B$, that we obtained in Proposition 3.6, by $\varphi_{B}$ and we will use $\varphi_{A}$ and $\psi_{A}$ for the left and the right integrals on the original discrete quantum group $A$. If no confusion is possible, we will drop these subscripts.
@@
    $a=(\iota\otimes\varphi)(\Delta(h)(a\otimes 1))=(\iota\otimes\varphi)(\Delta(h)(1\otimes S(a))$
    so that $\langle a,b\rangle=\varphi_{B}(b^{\prime}b)$ with $b^{\prime}=\varphi(\,\cdot\,S(a))$. Because $b\mapsto\varphi_{B}(b^{\prime}b)$ is continuous, this proves the result.
@@
  • Proof: Remark that $\pi(a)$ is defined as a multiplier of $\mathcal{K}(\mathcal{H})$ and so $\pi(a)$ is defined in $\mathcal{B}(\mathcal{H})$. The rest of the proof is essentially the same as for finite-dimensional corepresentations. $\square$
@@
Consider the GNS-representation of $B$ associated with the integral and denote by $\Lambda_{B}:B\to\mathcal{H}_{B}$ the associated canonical map. We let $B$ act directly on the Hibert space $\mathcal{H}_{B}$ so that $\Lambda_{B}(bb^{\prime})=b\Lambda_{B}(b^{\prime})$ for all $b,b^{\prime}\in B$. Similarly, take the GNS-representation of $A$ for the left integral $\varphi$ on $A$ and denote the canonical map by $\Lambda_{A}:A\to\mathcal{H}_{A}$.
@@
Finally, we look at some important formula we find in the theory of compact quantum groups and see what the corresponding results are for the discrete quantum groups.
@@
We also look at the dual of a compact group. In [VD-suq2] we treat a more advanced example in greater detail.
@@
Let us have a quick look at the formula in Proposition 2.11. For the element $q$, defined in Notation 2.10, we get here
@@
Indeed, the normalized trace on the components are given $\delta_{p}\mapsto 1$ for all $p$. So as expected, we get $q=1$ in this case.
@@
Compare the above argument with the proof of Proposition 3.6. There we get the positivity of the integral on $B$ from the positivity of the integral on $A$. Both results are essentially based on the same formula. This formula is like the Plancherel formula for locally compact abelian groups, see more details about this in [VD-part2].
@@
We are using that a multiplier Hopf -algebra with a positive integral is an operator algebra. In this case, because of the structure of $A$, this is an easy result. In general, it uses that the algebra is represented by bounded operator in the GNS representation of $A$ induced by the positive functional. See e.g. [VD-part2] for details.
@@
In this paper, we have given a treatment of discrete quantum groups. It is different from the original one in [VD-discrete]. The study of discrete quantum groups started before other more general developments in the theory of locally compact groups. We felt an update was needed with references to these newer developments.
An important role is played by the element $\Delta(h)$. It is a separability idempotent in the sense of [VD-si1] and many of its properties are in fact coming from this fact.
We consider the dual of a discrete quantum group in the sense of duality of algebraic quantum groups as in [VD-alg]. It is a compact quantum group and properties of the compact quantum group follow within this point of view of properties of discrete quantum groups.
In another paper on the subject [VD-suq2], the discrete quantum group of the quantized enveloping algebra of the Lie algebra of $SU(2)$ is studied in this setting. It illustrates many of the general results as described in this paper. It could be a challenging project to study more of such concrete examples in this framework.
@@
Compact quantum groups in the operator algebra framework are introduced and studied by Woronowicz. First he only considered a restricted class of such compact quantum groups in [Wo2] while later, he studied them from a more general point of view [Wo3]. This last paper only appeared in 1995, but it was available as a preprint many years before that. In this paper, we have used that a compact quantum group is in fact just a Hopf -algebra with a positive integral.
It seems appropriate to start this discussion with the work of Podleś and Woronowicz as they were the first to consider discrete quantum groups in an operator algebraic setting. We look at the Sections 2 and 3 in their paper on the quantum deformation of the Lorentz group [Po-Wo].
In Section 2, they first review the theory of compact matrix quantum groups as developed in [Wo2]. The underlying C-algebra is denoted by $A_{c}$ and we will denote the coproduct by $\Delta_{c}$. Then they use the results about corepresentations of $(A_{c},\Delta_{c})$ to construct the underlying algebra $A_{d}$ of the dual discrete quantum group. It is defined as a C-direct sum of matrix algebras. Corepresentation theory of $A_{c}$ is also used to construct the ’left regular corepresentation’ $u$ in the multiplier algebra $M(A_{d}\otimes A_{c})$ of the C-tensor product of $A_{d}$ with $A_{c}$.
The first main result in [Po-Wo] is Theorem 3.1. The coproduct on the dual $A_{d}$ is constructed, together with a counit and an antipode satisfying the standard equalities. The coproduct is defined as a non-degenerate -homomorphism $\Delta_{d}:A_{d}\to M(A_{d}\otimes A_{d})$ by the formula $(\Delta_{d}\otimes\iota)u=u_{23}u_{13}$. This means that they work with the ’flipped’ coproduct on the dual, i.e. the usual convention in the operator algebra approach. The counit is a -homomorphism from $A_{d}$ to $\mathbb{C}$ (and so everywhere defined and bounded). The antipode however is not defined everywhere but only on the algebraic direct sum, denoted by $A_{d0}$, of the components of $A_{d}$ and maps this dense -subalgebra into itself.
@@
So what is done by Podleś and Woronowicz in these two sections of [Po-Wo] can be summarized as follows. They define a discrete quantum group as a pair of a C-algebra AA with a coproduct $\Delta$ on $A$, where $A$ is a direct sum of matrix algebras and so that furthermore there is a counit and an antipode. They associate to each compact quantum matrix group a dual discrete quantum group and they obtain further results about this discrete quantum group (such as the existence of the integrals) using this fact and various results about compact quantum groups and their corepresentation theory.
@@
In [VD-discrete], a discrete quantum group is defined as a multiplier Hopf -algebra $(A,\Delta)$ where $A$ is a (algebraic) direct sum of matrix algebras (see Definition 2.3 in [VD-discrete]). A multiplier Hopf -algebra is defined in [VD-mha], Definitions 2.3 and 2.4.
This definition is not very different from the one given by Podleś and Woronowiz in the sense that it takes little effort to pass from one to the other. Indeed, if we start e.g. with a multiplier Hopf -algebra, the general theory allows to construct a counit and an antipode (see Section 3 and 4 in [VD-mha]). Then, as the algebra is a direct sum of matrix algebras, it is not hard to extend the coproduct and the counit to the C-direct sum. This takes us to a discrete quantum group as introduced (implicitly) in Theorem 3.1 of [Po-Wo].
Conversely, if we start with a discrete quantum group as in [Po-Wo], we simply restrict the coproduct to the algebraic direct sum of the components and we obtain a discrete quantum group in the sense of [VD-discrete]. The counit and the antipode are used to show that the basic axioms for a multiplier Hopf -algebra (see [VD-mha] and/or [VD-discrete]) are fulfilled. It is worthwhile noticing that the axioms for multiplier Hopf algebras can also be formulated in terms of the existence of a counit and an antipode (as for Hopf algebras) and then the two definitions come even closer together.

### v2

@@ Acknowledgments
  • Proof: i) We first look at the setting in [32]. For both the algebras $B$ and $C$ in that paper, we simply take $A$, and for $E$ we take $\Delta(h)$. The algebra $A$ is non-degenerate and $\Delta(h)$ is an idempotent in $M(A\otimes A)$. Here we have -algebras and $\Delta(h)$ is self-adjoint. We know that $\Delta(h)(1\otimes a)$ and $(a\otimes 1)\Delta(h)$ both belong to $A\otimes A$ and similarly on the other side. This shows that Assumption 1.1 of [32] are fulfilled. In particular $\Delta(h)$ is regular in the sense of that assumption.
@@
    that $S(e_{pj})\in V_{\overline{\alpha}}$ for all $j,p$. This implies that $V_{\overline{\alpha}}=A_{\overline{\alpha}}$ for all $\alpha$ and hence $V=A$. By the way, from Proposition 1.3 in [32] we can conclude that the legs of $\Delta(h)$ are all of $A$, now in the sense of our paper here. Also Proposition 1.4 in [32] here says that $\Delta(h)(1\otimes a)=0$ only if $a=0$, and similarly $(a\otimes 1)\Delta(h)=0$ only if $a=0$. This we knew already because we have a multiplier Hopf algebra and $h\neq 0$.
    iii) Finally we come to Definition 1.5 of [32]. Since we know here already that
    $\Delta(h)(a\otimes 1)=\Delta(h)(1\otimes S(a))\qquad\quad\text{and}\qquad\quad(1\otimes a)\Delta(h)=(S(a)\otimes 1)\Delta(h)$
@@
The maps $S:B\to C$ and $S^{\prime}:C\to B$ we have in Proposition 1.7 of [32] are precisely the antipode $S$ of $A$. If we look at the Proposition 1.8 in [32], we get here
@@
We consider the left and right integrals $\varphi$ and $\psi$, normalized such that $\varphi(h)=\psi(h)=1$. Then we have $\varphi\circ S=\psi$ as well as $\psi\circ S=\varphi$, see Proposition 1.12.
We take a positive invertible element in $M(A)$ satisfying $aq=S^{2}(a)q$ for all $a$. Then we get the following formulas for the integrals. Recall that we use $\omega_{\alpha}$ for the standard trace on $A_{\alpha}$, see Notation 2.15.
@@
  • Theorem When $a\in A_{\alpha}$ we have
    $\varphi(a)=\omega_{\alpha}(q)\omega_{\alpha}(aq^{-1}).$ (2.4)
  • Proof: Define $\varphi$ on $A$ by $\varphi(a)=\omega_{\alpha}(q)\omega_{\alpha}(aq^{-1})$ when $a\in A_{\alpha}$. We have to show that $(\iota\otimes\varphi)\Delta(h)=1$. This means that we need $(\iota\otimes\varphi)(\Delta(h)(1_{\overline{\alpha}}\otimes 1))=1_{\overline{\alpha}}$ for all $\alpha$.
    Now $\Delta(h)(1_{\overline{\alpha}}\otimes 1)=\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha})$ and we can use the formula of Proposition 2.14. We get
    $(\iota\otimes\varphi)(\Delta(h)(1_{\overline{\alpha}}\otimes 1_{\alpha}))$ $=\frac{1}{c}\sum_{i,j}\lambda_{i}\,S(e_{ij})\varphi(e_{ji})$
    $=\frac{1}{c}\sum_{i,j}\lambda_{i}\,S(e_{ij})c\omega_{\alpha}(e_{ji}q^{-1})$
@@
    $=\sum_{i}S(e_{ii})=S(1_{\alpha})=1_{\overline{\alpha}}.$
    This proves the result.
Observe that the right hand side of Equation (2.4) does not depend on the choice of $q$.
@@
  • Theorem For any $a\in A_{\alpha}$ we have
    $\psi(a)=\omega_{\alpha}(q^{-1})\omega_{\alpha}(aq).$ (2.5)
  • Proof: i) Take $a\in A_{\overline{\alpha}}$. Then $S^{-1}(a)\in A_{\alpha}$ and we can apply the formula for $\varphi$. We find
    $\psi(a)=\varphi(S^{-1}(a))$ $=\omega_{\alpha}(q)\omega_{\alpha}(S^{-1}(a)q^{-1})$
    $=\omega_{\alpha}(q)\omega_{\alpha}(S^{-1}(S(q^{-1})a))$
    $=\omega_{\alpha}(q)\omega_{\overline{\alpha}}(S(q^{-1})a)$
    $=\omega_{\alpha}(q)\omega_{\overline{\alpha}}(aS(q^{-1})).$
    ii) Denote $q^{\prime}=S(q^{-1})$ so that $q=S^{-1}({q^{\prime}}^{-1})$ and $\omega_{\alpha}(q)=\omega_{\overline{\alpha}}({q^{\prime}}^{-1})$. If we apply $S$ on the equation $aq=qS^{2}(a)$ we get $S(q)S(a)=S^{2}(S(a))S(q)$. This holds for all $a$ and because $S$ is bijective we also have $S(q)a=S^{2}(a)S(q)$ and $aS(q^{-1})=S(q^{-1})S^{2}(a)$. As $q^{\prime}=S(q^{-1})$ we get $aq^{\prime}=q^{\prime}S^{2}(a)$ for all $a$.
    iii) We see that $\psi(a)=\omega_{\overline{\alpha}}({q^{\prime}}^{-1})\omega_{\overline{\alpha}}(aq^{\prime})$ for $a\in A_{\overline{\alpha}}$. We can scale $q^{\prime}$ to get a positive invertible element. Then the result follows for $\overline{\alpha}$ instead of $\alpha$.
Also here, the right hand side of Equation (2.5) does not depend on the choice of $a$.
We now obtain formulas for the other data, the modular automorphism $\sigma$, the modular element $\delta$. We also verify some formulas relating these data.
2.20.
  • Proposition The $\alpha$-component $\delta_{\alpha}$ of the modular element $\delta$ is given by
    $\delta_{\alpha}=\frac{\omega_{\alpha}(q^{-1})}{\omega_{\alpha}(q)}\,q_{\alpha}^{2}.$ (2.6)
  • Proof: The results follows from the two previous ones because $\psi(a)=\varphi(a\delta)$ for all $a$ in $A$.
The right hand side of equality Equation (2.6 does not depend on the choice of $q$ because is we replace $q_{\alpha}$ by a multiple of it, we get the same result.
@@
We can also verify the formulas
$(\varphi\otimes\iota)\Delta(h)=\delta\qquad\quad\text{and}\qquad\quad(\iota\otimes\psi)\Delta(h)=\delta^{-1}.$
For the first one we have
$(\varphi\otimes\iota)(\Delta(h)(1\otimes 1_{\alpha}))$ $=\frac{1}{\omega_{\alpha}(q)}\sum_{i,j}\lambda_{i}\varphi(S(e_{ij}))e_{ji}$
$=\frac{1}{\omega_{\alpha}(q)}\sum_{i,j}\lambda_{i}\psi(e_{ij})e_{ji}$
$=\frac{1}{\omega_{\alpha}(q)}\sum_{i}\omega_{\alpha}(q^{-1})\lambda_{i}^{2}e_{ii}$
$=\frac{\omega_{\alpha}(q^{-1})}{\omega_{\alpha}(q)}\sum_{i}\lambda_{i}^{2}e_{ii}$
$=\frac{\omega_{\alpha}(q^{-1})}{\omega_{\alpha}(q)}q_{\alpha}^{2}=\delta_{\alpha}.$
For the second equation we have
$(\iota\otimes\psi)(\Delta(h)(1_{\overline{\alpha}}\otimes 1))$ $=\frac{1}{\omega_{\alpha}(q)}\sum_{i,j}\lambda_{i}S(e_{(ij})\psi(e_{ji}$
$=\frac{1}{\omega_{\alpha}(q)}\sum_{i}\lambda_{i}S(e_{(ii})\omega_{\alpha}(q^{-1})\lambda_{i}$
$=\frac{\omega_{\alpha}(q^{-1})}{\omega_{\alpha}(q)}S(q_{\alpha}^{2})=S(\delta_{\alpha})=\delta^{-1}_{\overline{\alpha}}.$
@@
$\varphi(ab)=\omega_{\alpha}(q)\omega_{\alpha}(abq_{\alpha}^{-1})=\omega_{\alpha}(q)\omega_{\alpha}(bq_{\alpha}^{-1}a)=\omega_{\alpha}(q)\varphi(bq_{\alpha}^{-1}aq_{\alpha}).$
We see that $\sigma(a)=q_{\alpha}^{-1}aq_{\alpha}$ and this is in agreement with $\sigma(a)=S^{2}(a)$.
Similarly, from $\psi(a)=\omega_{\alpha}(q^{-1})\omega_{\alpha}(aq)$ when $a\in A_{\alpha}$, we get $\sigma^{\prime}(a)=q_{\alpha}aq_{\alpha}^{-1}$. This is in agreement with $\sigma^{\prime}(a)=S^{-2}(a).$
@@
In this section, we will consider the dual of a discrete quantum group. We will follow the approach to duality as developed for general algebraic quantum groups in [29]. The dual will be a compact quantum group and we know that all compact quantum groups arise in this way. By looking at a compact quantum group as the dual of a discrete quantum group, we can obtain results about compact quantum groups in a different and easier way. We refer to the appendix for more comments on this matter.
@@
    iii) In this case, we also use $\Delta$ for the coproduct, $\varepsilon$ for the counit, $S$ for the antipode on $B$.
@@
Again observe the presence of the inverse of the antipode in Equation (3.3). This is one of the consequences of this convention. For the same reason, there is the lack of symmetry between (3.4) and (3.5). We do not get the usual formulas for a pairing of Hopf -algebras or multiplier Hopf -algebras (as e.g. in [3]), but rather a modified form.
@@
So we see that the dual of the discrete quantum group (seen as a multiplier Hopf -algebra with positive integrals) is a Hopf -algebra with positive integrals. We also see from the above equation that the integral is faithful. This is in fact automatic (see e.g. [29]). We will encounter a similar result with a similar argument in Proposition 3.13 later in this section, as well in Proposition 4.6 in the next section with a dual form.
@@
  • Proof: i) Let $d=\varphi_{A}(a\,\cdot\,)$ with $a\in A$. Then for all $c\in B$,
@@
    $=\sum_{(h)}\langle h_{(1)},c\rangle\varphi_{A}(ah_{(2)})$
    $=\sum_{(h)}\langle S(a)h_{(1)},c\rangle\varphi_{A}(h_{(2)})$
@@
    $=\sum_{(h)}\varphi_{A}(ah_{(1)})\langle h_{(2)},b\rangle)$
    $=\sum_{(h)}\varphi_{A}(h_{(1)})\langle S^{-1}(a)h_{(2)},b\rangle)$
@@
Formulas of this type are found at various places in the literature. One possible reference is [2] where they are obtained in the more general case of algebraic quantum hypergroups. They are also found in [35].
If we let $B$ act on the GNS-representation space associated with this integral, we can complete it to a C-algebra. Also the coproduct can be extended. One of the key features that makes this possible is the fact that the -algebra is represented by bounded operators in the GNS-space associated with the positive integral. This is not true for general positive linear functionals on a -algebra. This result is already found in [13], see also [1] and the more recent approach in [36].
In what follows we will mostly denote the integral on the dual compact quantum group $B$, that we obtained in Proposition 3.6, by $\varphi_{B}$ and we will use $\varphi_{A}$ and $\psi_{A}$ for the left and the right integrals on the original discrete quantum group $A$. If no confusion is possible, we will drop these subscripts.
@@
    $a=(\iota\otimes\varphi_{A})(\Delta(h)(a\otimes 1))=(\iota\otimes\varphi_{A})(\Delta(h)(1\otimes S(a))$
    so that $\langle a,b\rangle=\varphi_{B}(b^{\prime}b)$ with $b^{\prime}=\varphi_{A}(\,\cdot\,S(a))$. Because $b\mapsto\varphi_{B}(b^{\prime}b)$ is continuous, this proves the result.
@@
  • Proof: Remark that $\pi(a)$ is defined as a multiplier of $\mathcal{K}(\mathcal{H})$ and so $\pi(a)$ is defined in $\mathcal{B}(\mathcal{H})$. The rest of the proof is essentially the same as for finite-dimensional corepresentations.
@@
Consider the GNS-representation of $B$ associated with the integral and denote by $\Lambda_{B}:B\to\mathcal{H}_{B}$ the associated canonical map. We let $B$ act directly on the Hibert space $\mathcal{H}_{B}$ so that $\Lambda_{B}(bb^{\prime})=b\Lambda_{B}(b^{\prime})$ for all $b,b^{\prime}\in B$. Similarly, take the GNS-representation of $A$ for the left integral $\varphi_{A}$ on $A$ and denote the canonical map by $\Lambda_{A}:A\to\mathcal{H}_{A}$.
@@
Finally, we look at some important formulas we find in the theory of compact quantum groups and see what the corresponding results are for the discrete quantum groups.
@@
We also look at the dual of a compact group. In [41] we treat a more advanced example in greater detail.
@@
Let us have a quick look at the formula in Proposition 2.9. For the standard trace on the components, we simply have $\omega(\delta_{p})=1$ for all $p$. For the element $q$, defined in Proposition 2.9, we get here
@@
Because we took the normalized trace, we get $q=1$ in this case as expected.
@@
Compare the above argument with the proof of Proposition 3.6. There we get the positivity of the integral on $B$ from the positivity of the integral on $A$. Both results are essentially based on the same formula. This formula is like the Plancherel formula for locally compact abelian groups, see more details about this in [36].
@@
We are using that a multiplier Hopf -algebra with a positive integral is an operator algebra. In this case, because of the structure of $A$, this is an easy result. In general, it uses that the algebra is represented by bounded operator in the GNS-representation of $A$ induced by the positive functional. See e.g. [36] for details.
@@
In this paper, we have given a treatment of discrete quantum groups. It is different from the original one in [28]. The study of discrete quantum groups started before other more general developments in the theory of locally compact groups. We felt an update was needed with references to these newer developments.
An important role is played by the element $\Delta(h)$. It is a separability idempotent in the sense of [32] and many of its properties are in fact coming from this fact.
We consider the dual of a discrete quantum group in the sense of duality of algebraic quantum groups as in [29]. It is a compact quantum group and properties of the compact quantum group follow within this point of view of properties of discrete quantum groups.
In another paper on the subject [41], the discrete quantum group of the quantized enveloping algebra of the Lie algebra of $SU(2)$ is studied in this setting. It illustrates many of the general results as described in this paper. It could be a challenging project to study more of such concrete examples in this framework.
@@
Compact quantum groups in the operator algebra framework are introduced and studied by Woronowicz. First he only considered a restricted class of such compact quantum groups in [42] while later, he studied them from a more general point of view [44]. This last paper only appeared in 1995, but it was available as a preprint many years before that. In this paper, we have used that a compact quantum group is in fact just a Hopf -algebra with a positive integral.
It seems appropriate to start this discussion with the work of Podleś and Woronowicz as they were the first to consider discrete quantum groups in an operator algebraic setting. We look at the Sections 2 and 3 in their paper on the quantum deformation of the Lorentz group [18].
In Section 2, they first review the theory of compact matrix quantum groups as developed in [42]. The underlying C-algebra is denoted by $A_{c}$ and we will denote the coproduct by $\Delta_{c}$. Then they use the results about corepresentations of $(A_{c},\Delta_{c})$ to construct the underlying algebra $A_{d}$ of the dual discrete quantum group. It is defined as a C-direct sum of matrix algebras. Corepresentation theory of $A_{c}$ is also used to construct the ’left regular corepresentation’ $u$ in the multiplier algebra $M(A_{d}\otimes A_{c})$ of the C-tensor product of $A_{d}$ with $A_{c}$.
The first main result in [18] is Theorem 3.1. The coproduct on the dual $A_{d}$ is constructed, together with a counit and an antipode satisfying the standard equalities. The coproduct is defined as a non-degenerate -homomorphism $\Delta_{d}:A_{d}\to M(A_{d}\otimes A_{d})$ by the formula $(\Delta_{d}\otimes\iota)u=u_{23}u_{13}$. This means that they work with the ’flipped’ coproduct on the dual, i.e. the usual convention in the operator algebra approach. The counit is a -homomorphism from $A_{d}$ to $\mathbb{C}$ (and so everywhere defined and bounded). The antipode however is not defined everywhere but only on the algebraic direct sum, denoted by $A_{d0}$, of the components of $A_{d}$ and maps this dense -subalgebra into itself.
@@
So what is done by Podleś and Woronowicz in these two sections of [18] can be summarized as follows. They define a discrete quantum group as a pair of a C-algebra AA with a coproduct $\Delta$ on $A$, where $A$ is a direct sum of matrix algebras and so that furthermore there is a counit and an antipode. They associate to each compact quantum matrix group a dual discrete quantum group and they obtain further results about this discrete quantum group (such as the existence of the integrals) using this fact and various results about compact quantum groups and their corepresentation theory.
@@
In [28], a discrete quantum group is defined as a multiplier Hopf -algebra $(A,\Delta)$ where $A$ is a (algebraic) direct sum of matrix algebras (see Definition 2.3 in [28]). A multiplier Hopf -algebra is defined in [27], Definitions 2.3 and 2.4.
This definition is not very different from the one given by Podleś and Woronowiz in the sense that it takes little effort to pass from one to the other. Indeed, if we start e.g. with a multiplier Hopf -algebra, the general theory allows to construct a counit and an antipode (see Section 3 and 4 in [27]). Then, as the algebra is a direct sum of matrix algebras, it is not hard to extend the coproduct and the counit to the C-direct sum. This takes us to a discrete quantum group as introduced (implicitly) in Theorem 3.1 of [18].
Conversely, if we start with a discrete quantum group as in [18], we simply restrict the coproduct to the algebraic direct sum of the components and we obtain a discrete quantum group in the sense of [28]. The counit and the antipode are used to show that the basic axioms for a multiplier Hopf -algebra (see [27] and/or [28]) are fulfilled. It is worthwhile noticing that the axioms for multiplier Hopf algebras can also be formulated in terms of the existence of a counit and an antipode (as for Hopf algebras) and then the two definitions come even closer together.